### v1

Abstract
Nonsmooth composite optimization problems under uncertainty are prevalent in various scientific and engineering applications. We consider risk-neutral composite optimal control problems, where the objective function is the sum of a potentially nonconvex expectation function and a nonsmooth convex function. To approximate the risk-neutral optimization problems, we use a Monte Carlo sample-based approach, study its asymptotic consistency, and derive nonasymptotic sample size estimates. Our analyses leverage problem structure commonly encountered in PDE-constrained optimization problems, including compact embeddings and growth conditions. We apply our findings to bang-bang-type optimal control problems and propose the use of a conditional gradient method to solve them effectively. We present numerical illustrations.
Key words. bilinear control, sample average approximation, empirical risk minimization, PDE-constrained optimization, stochastic optimization, bang-bang control
AMS subject classifications. 90C15, 90C59, 90C06, 35Q93, 49M41, 90C60
1 Introduction
In this paper, we consider composite risk-neutral optimization problems of the form
$\min_{u\in U}\,G(u)\coloneqq F(Bu)+\psi(u),\quad\text{where}\quad F(w)\coloneqq\mathbb{E}[f(w,\boldsymbol{\xi})],$ ($\mathcal{P}$)
$B\colon U\to W$ is a compact linear operator between the Hilbert space $U$ and the Banach space $W$, $f\colon W\times\Xi\to[0,\infty)$ is a smooth integrand, and $\psi:U\to[0,\infty]$ is a proper, closed, and convex but potentially nonsmooth. Moreover, $\boldsymbol{\xi}$ denotes a random element, which maps to a complete, separable metric space $\Xi$. While ($\mathcal{P}$) covers a variety of challenging settings, the focus of the present work lies on risk-neutral PDE-constrained optimal control problems
$\min_{u\in L^{2}(D)}\,\mathbb{E}[J(S(Bu,\boldsymbol{\xi}))]+\psi(u),$ (1.1)
where $J$ is a tracking-type functional, and the choice of the deterministic control $u$ influences the behavior of the solution $S(Bu,\boldsymbol{\xi})$ to a PDE with random inputs $\boldsymbol{\xi}$.
Nonconvex optimization problems governed by differential equations arise in a multitude of application areas, such as sensor placement [2], resource assessment of renewable tidal energy [8], and design of groundwater remediation systems [9]. It is well known that the proper choice of the penalty function $\psi$ in Problem 1.1 promotes certain structural features in its minimizers $u^{*}$. For example, setting
$\psi(u)=I_{U_{\text{ad}}}(u)+\beta\|u\|_{L^{1}(D)},\quad\text{where}\quad U_{\text{ad}}\coloneqq\left\{\,u\in L^{2}(D)\;|\;-1\leq u(x)\leq 1\quad\text{for a.e.}~{}x\in D\right\},$
$I_{U_{\text{ad}}}$ denotes its indicator function, and $\beta>0$, tends to provide solutions satisfying a bang-bang-off principle, that is, $u^{*}(x)\in\{-1,0,1\}$ for a.e. $x\in D$. As a consequence, composite optimal control problems also arise from relaxations of mixed-integer optimal control problems [8].
In order to handle the expected value in ($\mathcal{P}$) numerically, we approximate it using the sample average approximation (SAA) approach [12]. More specifically, given a sequence $\boldsymbol{\xi}^{1},\boldsymbol{\xi}^{2},\ldots,\boldsymbol{\xi}^{N},\ldots$ defined on a complete probability space $(\Omega,\mathcal{F},P)$ of independent $\Xi$-valued random elements such that each $\boldsymbol{\xi}^{i}$ has the same distribution as $\boldsymbol{\xi}$, we consider the SAA problem
$\min_{u\in U}\,\hat{F}_{N}(Bu,\omega)+\psi(u),\quad\text{where}\quad\hat{F}_{N}(w,\omega)\coloneqq\frac{1}{N}\sum_{i=1}^{N}f(w,\boldsymbol{\xi}^{i}(\omega)).$ ($\mathcal{P}_{N}$)
For brevity, we often omit the second argument of $\hat{F}_{N}$.
A central topic arising in this context is the question whether this type of approximation is asymptotically consistent and if the proximity of its solutions and critical points towards their counterpart in ($\mathcal{P}$) can be quantified in terms of the sample size $N$. Although this already represents a formidable problem for smooth objective functions, the potential lack of strong convexity of $\psi$ aggravates the problem further, adding additional challenges to the theoretical analysis of the problem and its efficient solution.
1.1 Contributions
In the present paper, we investigate the SAA approach for nonsmooth, potentially nonconvex risk-neutral minimization problems of the form ($\mathcal{P}$) from a qualitative and a quantitative perspective. Since Problem ($\mathcal{P}$) is potentially nonconvex, our analysis addresses the behaviour its solutions and its critical points. In the context of the present manuscript, the latter refers to points $u^{*}_{N}\in U$ satisfying the first-order necessary subdifferential inclusion
$-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*}_{N})\in\partial\psi(u^{*}_{N})$
or, equivalently, zeros of the gap functional
$\hat{\Psi}_{N}(u)$ $\coloneqq\sup_{v\in U}\,[\,(B^{*}\mathrm{D}\hat{F}_{N}(Bu),u-v)_{U}+\psi(u)-\psi(v)\,].$ (1.2)
Critical points as well as the associated gap function $\Psi$ for Problem ($\mathcal{P}$) are defined analogously. More specifically, the contributions of the paper are as follows:
  1. (a)
    For general smooth fidelity terms $F$, we show that the random variables realizing the optimal values, the solutions, and the critical points of Problem ($\mathcal{P}_{N}$) asymptotically approximate their deterministic counterparts of Problem ($\mathcal{P}$) with probability one (w.p. $1$) as the sample size $N$ tends to infinity.
  2. (b)
    We establish nonasymptotic expectation bounds for the gap functional value of arbitrary sequences $(u_{N})$ of random vectors. More precisely, we demonstrate that, for every $\varepsilon>0$ there holds
    $\mathbb{E}[\Psi(u_{N})]\leq\mathbb{E}[\hat{\Psi}_{N}(u_{N})]+\varepsilon\quad\text{if}\quad N\geq\frac{c_{1}}{\varepsilon^{2}}\cdot\mathcal{N}(\varepsilon/c_{1};B(\mathcal{D}(\psi))),$
    where $\mathcal{N}(\nu;B(\mathcal{D}(\psi)))$, $\nu>0$, denotes the $\nu$-covering number of the compact set $B(\mathcal{D}(\psi))\subset W$ and $c_{1}>0$ explicitly depends on the radius of the domain of $\psi$ as well as the smoothness of the integrand $f$.
  3. (c)
    While Problem ($\mathcal{P}$) is typically not strongly convex, we provide nonaymptotic expectation bounds for the approximation of minimizers $u^{*}$ to ($\mathcal{P}$) by random vectors $(u^{*}_{N})$ realizing solutions of Problem ($\mathcal{P}_{N}$) if the integrand is convex for each $\xi\in\Xi$, and a growth-type condition on the partial linearization of $G$ at a minimizer $u^{*}$ holds true. Specifically, if
    $(B^{*}\mathrm{D}F(Bu^{*}),u-u^{*})_{U}+\psi(u)-\psi(u^{*})\geq\theta\|u-u^{*}\|_{\mathcal{U}}^{2}\quad\text{for all}\quad u\in U$ (1.3)
    for some $\theta>0$ and a Banach space $(\mathcal{U},\|\cdot\|_{\mathcal{U}})$ with $U\hookrightarrow\mathcal{U}$, then
    $\mathbb{E}[\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}^{2}]+\mathbb{E}[\Psi(u^{*}_{N})]+\mathbb{E}[G(u^{*}_{N})-G(u^{*})]\leq(1/N)c_{2},$
    where $c_{2}>0$ depends on the smoothness of the integrand, and $\theta$. Hence, the SAA solutions converge in expectation at the usual Monte Carlo rate of $(1/N)^{1/2}$ while both the value of the gap functional and the suboptimality in $G$ exhibit superconvergence at a rate of $(1/N)$. This convergence statement is valid for linear bounded operators $B$ without their compactness.
Our framework is applied to PDE-constrained optimization problems governed by affine-linear and bilinear elliptic equations, which allow for the use of bang-bang-off regularization terms. Reproducible numerical experiments empirically verify our theoretical results and further highlight the utility of the SAA approach for infinite-dimensional, nonsmooth problems.
1.2 Related work
Monte Carlo sample-based approximations are common discretization approaches for risk-neutral and risk-averse huge-scale optimization, particularly in PDE-constrained optimization [19], and stochastic optimization [34]. The theoretical analyses of this approximation approach may be categorized into asymptotic and nonasymptotic ones. In PDE-constrained optimization under uncertainty with strongly convex control regularizers, the papers [11, 18, 20, 31] provide nonasymptotic analyses, such as sample size estimates and central limit-type theorems. Moreover, [19, 24, 25] provide asymptotic consistency results for nonconvex infinite-dimensional stochastic optimization. The problem structure given by the compact linear operator in (4.1) is common among PDE-constrained optimization problems and has been used for different purposes in the literature. For example, the authors of [25] have employed it to demonstrate the consistency of empirical estimators for risk-averse stochastic optimization. Finally, [26] establishes nonasymptotic sample size estimates for risk-neutral semilinear PDE-constrained optimization. In the field of PDE-constrained optimization under uncertainty, current SAA analyzes require certain control regularizers, such as standard Tikhonov regularizers [11, 21], R-functions [25], and Kadec functions [24]. These requirements exclude the $L^{1}$-norm as a control regularizer, for example. Using the problem structure given by the compact linear operator in (4.1), we are able to extend previous results to general convex control regularizers.
As mentioned in the previous section, our sample size estimates for convex problems do not depend on the compactness of the feasible set or of the operator $B$. Sample size estimates for stochastic convex optimization without any compactness conditions are also established, for example, in [15, 21, 33].
1.3 Outline
Section 2 introduces notation and terminology. Section 3 introduces assumptions on the risk-neutral composite control problems, studies the existence of solutions, states gap functional-based optimality conditions, and establishes consistency of SAA optimal values, solutions and critical points. Section 3 also derives sample size estimates for nonconvex and convex problems. Our framework is applied to linear and bilinear PDE-constrained optimization in section 4. Section 5 presents numerical illustrations. The appendix derives uniform expectation bounds.
2 Notation and terminology
Throughout the text, normed vector spaces $X$ are defined over the reals. Metric spaces $X$ are equipped with their Borel sigma-field $\mathcal{B}(X)$. Moreover, we identify the control space $U$, a separable Hilbert space, with $U^{*}$, and write $U^{*}=U$. For a Banach space $(X,\|\cdot\|_{X})$ with norm $\|\cdot\|_{X}$, we denote by $X^{*}$ its topological dual space and by $\langle\cdot,\cdot\rangle_{{X}^{*},X}$ its duality pairing. If the norm is clear from the context, we write $X$ instead of $(X,\|\cdot\|_{X})$. The inner product on a Hilbert space $H$ is denoted by $(\cdot,\cdot)_{H}$. If $X$ and $Y$ are Banach spaces with $X\subset Y$, and $\imath\colon X\to Y$ defined by $\imath x\coloneqq x$ is linear and bounded, then we say that $X$ is (continuously) embedded into $Y$. We abbreviate such embeddings by $X\hookrightarrow Y$. A linear operator between two Banach spaces is called compact if its image of the domain space’s unit ball is precompact. For a linear bounded operator $\Upsilon$ between Banach spaces, $\Upsilon^{*}$ denotes its adjoint operator. For a measurable space $(\Theta,\mathcal{A})$ and metric spaces $X$ and $Y$, $h\colon X\times\Theta\to Y$ is called Carathéodory mapping if $h(\cdot,\theta)$ is continuous for all $\theta\in\Theta$ and $h(x,\cdot)$ is $\mathcal{A}$-$\mathcal{B}(Y)$ measurable for all $x\in X$. For a bounded domain $D\subset\mathbb{R}^{d}$, we denote by $L^{p}(D)$ $(1\leq p\leq\infty)$ the standard Lebesgue spaces, and by $H_{0}^{1}(D)$ and $H^{s}(D)$ ($s\in\mathbb{N})$ the standard Sobolev spaces. We set $H^{-1}(D)\coloneqq H_{0}^{1}(D)^{*}$. For a nonempty, totally bounded metric space $Y$, the $\nu$-covering number $\mathcal{N}(\nu;Y)$, $\nu>0$, denotes the minimal number of points in a $\nu$-net of $Y$.
3 Risk-neutral composite optimal control
We start by stating the main assumptions on Problem ($\mathcal{P}$). For this purpose, we abbreviate the domain of $\psi$ by
$\mathcal{D}(\psi)\coloneqq\{\,u\in U\;|\;\psi(u)<\infty\,\}.$
The following assumption imposes the properties of $\psi$ and the smoothness and integrability requirements on the integrand.
Assumption 1 (Integrand and feasible set).
  1. (a)
    The control space $U$ is a separable Hilbert space, and $W$ is a separable Banach space.
  2. (b)
    The operator $B:U\to W$ is linear, and compact.
  3. (c)
    The function $\psi:U\to[0,\infty]$ is proper, closed, convex, and $\mathcal{D}(\psi)$ is bounded.
  4. (d)
    The set $U_{0}\subset U$ is open, convex, and bounded with $\mathcal{D}(\psi)\subset U_{0}$. Moreover, $f\colon B(U_{0})\times\Xi\to\mathbb{R}$ is continuous in its first argument on $B(\mathcal{D}(\psi))$ for each $\xi\in\Xi$ and measurable in its second argument on $\Xi$ for each $u\in B(U_{0})$.
  5. (e)
    For an integrable random variable $\zeta_{f}:\Xi\to[0,\infty)$, it holds that
    $|f(Bu,\xi)|\leq\zeta_{f}(\xi)\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi.$
The following assumption formulates differentiablity and integrability statements of derivatives.
Assumption 2 (Integrand and its gradient).
  1. (a)
    For each $\xi\in\Xi$, the mapping
    $g_{\xi}\colon U_{0}\to\mathbb{R},\quad\text{where}\quad g_{\xi}(u)\coloneqq f(Bu,\xi),$
    is continuously differentiable.
  2. (b)
    There exists a Carathéodory function $\mathrm{D}_{w}f\colon B(\mathcal{D}(\psi))\times\Xi\to W^{*}$ such that
    $\nabla g_{\xi}(u)=B^{*}\mathrm{D}_{w}f(Bu,\xi)\quad\text{for all}\quad(u,\xi)\in\mathcal{D}(\psi)\times\Xi.$
  3. (c)
    There exists an integrable random variable $\zeta_{\mathrm{D}f}:\Xi\to[0,\infty)$ such that
    $\|\nabla g_{\xi}(u)\|_{U}$ $\leq\zeta_{\mathrm{D}f}(\xi)\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi,\quad\quad\text{and}$
    $\|\mathrm{D}_{w}f(Bu,\xi)\|_{W^{*}}$ $\leq\zeta_{\mathrm{D}f}(\xi)\quad\text{for all}\quad(u,\xi)\in\mathcal{D}(\psi)\times\Xi.$
If $f(\cdot,\xi)$ is continuously differentiable on a $W$-neighborhood of $B(\mathcal{D}(\psi))$, the differentiability of $g_{\xi}(\cdot)=f(B\cdot,\xi)$ and the representation of its gradient are implied by the chain rule. From this perspective, Assumption 2 will ensure that the gradients of $F\circ B$ and $\hat{F}_{N}\circ B$ admit a composite structure even in the absence of the chain rule. While this might seem technical at first glance, it will allow us to fit challenging settings into the outlined abstract framework. For a particular example, we refer the reader to the bilinear PDE-constrained problem in section 4.2.
3.1 Existence of solutions and necessary optimality conditions
In this section, we show that both the risk-neutral problem ($\mathcal{P}$) and the associated SAA problems ($\mathcal{P}_{N}$) admit solutions. We also show the measurability of the SAA optimal value and the existence of measurable SAA solutions. Moreover, we introduce the particular form of first-order necessary optimality conditions used in the following sections.
For $\omega\in\Omega$ and $N\in\mathbb{N}$, we denote by $\hat{\vartheta}^{*}_{N}(\omega)$ the optimal value of the SAA problem ($\mathcal{P}_{N}$).
Proposition 3.1.
Let Assumption 1 hold, and let $N\in\mathbb{N}$. Then, (i) the risk-neutral problem ($\mathcal{P}$) and for each $\omega\in\Omega$, the SAA problem ($\mathcal{P}_{N}$) admit solutions, (ii) the function $\hat{\vartheta}^{*}_{N}\colon\Omega\to\mathbb{R}$ is measurable, and (iii) there exists at least one measurable map $u^{*}_{N}\colon\Omega\to U$ such that for each $\omega\in\Omega$, $u^{*}_{N}(\omega)$ solves the SAA problem ($\mathcal{P}_{N}$). (iv) If, moreover, LABEL: and 2 holds, then the composite functions $u\mapsto F(Bu)$ and $u\mapsto\hat{F}_{N}(Bu)$ are continuously differentiable on $U_{0}$ with gradients $u\mapsto\mathbb{E}[\nabla g_{\boldsymbol{\xi}}(u)]$ and $u\mapsto(1/N)\sum_{i=1}^{N}\nabla g_{\boldsymbol{\xi}^{i}}(u)$, respectively.
Proof.
(i) Since $\mathcal{D}(\psi)$ is weakly compact, the proof follows by standard arguments as well as Fatou’s lemma.
(ii)–(iii) Since $(u,t,\omega)\mapsto\hat{F}_{N}(Bu,\omega)+t$ is Carathéodory on $\{(u,t)\in\mathcal{D}(\psi)\times\mathbb{R}\colon\psi(u)\leq t\}\times\Omega$, measurability theorems on marginal maps and inverse images [4, Thms. 8.2.9 and 8.2.11] imply the assertions.
(iv) The assertions can be established using standard arguments. ∎
Propositions 3.1 and 2 motivate the definitions
$\mathrm{D}F(w)\coloneqq\mathbb{E}[\mathrm{D}_{w}f(w,\boldsymbol{\xi})]\quad\text{and}\quad\mathrm{D}\hat{F}_{N}(w,\omega)\coloneqq\frac{1}{N}\sum_{i=1}^{N}\mathrm{D}_{w}f(w,\boldsymbol{\xi}^{i}(\omega))$
as mappings on $B(\mathcal{D}(\psi))$ and $B(\mathcal{D}(\psi))\times\Omega$, respectively. We note that these definitions are formal in the sense that while $\mathbb{E}[\mathrm{D}_{w}f(w,\boldsymbol{\xi})]$ is well-defined, $F$ may not be continuously differentiable, see also the discussion following Assumption 2. As with $\hat{F}_{N}$, we often omit the second argument of $\mathrm{D}\hat{F}_{N}$.
Furthermore, we define the gap functionals $\Psi:\mathcal{D}(\psi)\to[0,\infty]$ and $\hat{\Psi}_{N}\colon\mathcal{D}(\psi)\to[0,\infty]$ by
[…]
A point $\bar{u}\in\mathcal{D}(\psi)$ is referred to as critical point of ($\mathcal{P}$) if $-B^{*}\mathrm{D}F(B\bar{u})\in\partial\psi(\bar{u})$. Similarly, a point $\bar{u}_{N}\in\mathcal{D}(\psi)$ is called a critical point of the SAA problem ($\mathcal{P}_{N}$) if $-B^{*}\mathrm{D}\hat{F}_{N}(\bar{u}_{N})\in\partial\psi(\bar{u}_{N})$. Critical points of ($\mathcal{P}$) are the zeros of the gap function $\Psi$ as summarized in the following proposition.
Proposition 3.2.
If Assumptions 1 and 2 hold, then (i) an element $\bar{u}\in\mathcal{D}(\psi)$ is a critical point of ($\mathcal{P}$) if and only if $\Psi(\bar{u})=0$, (ii) a point $\bar{u}_{N}\in\mathcal{D}(\psi)$ is a critical point of ($\mathcal{P}_{N}$) if and only if $\hat{\Psi}_{N}(\bar{u}_{N})=0$, and (iii) $\Psi$ is weakly lower semicontinuous.
Proof.
Parts (i) and (ii) can be established using arguments similar to those used to prove Theorem 2.4 in [13]. Part (iii) follows from Theorem 2.5 (c) in [10] because $B$ is completely continuous. ∎
3.2 Consistency of SAA optimal values and solutions
We establish the asymptotic consistency of SAA optimal values and the weak consistency of SAA solutions. We denote by $\vartheta^{*}$ the optimal value of the true problem ($\mathcal{P}$).
Theorem 3.3.
If Assumption 1 holds, then (i) $\hat{\vartheta}_{N}^{*}\to\vartheta^{*}$ as $N\to\infty$ w.p. $1$, and (ii) for almost all $\omega\in\Omega$, $(u^{*}_{N}(\omega))$ has at least one weak accumulation point in $U$ and every such point is a solution to ($\mathcal{P}$).
Proof.
Before we establish parts (i) and (ii), let us note that the uniform law of large numbers [14, Cor. 4:1] ensures the existence of a null set $\Omega_{0}\subset\Omega$ with $\Omega_{0}\in\mathcal{F}$ such that for all $\omega\in\Omega\setminus\Omega_{0}$, we have
$\sup_{w\in B(\mathcal{D}(\psi))}|F(w)-\hat{F}_{N}(w,\omega)|\to 0\quad\text{as}\quad N\to\infty.$ (3.2)
Moreover, $F$ is continuous on $B(\mathcal{D}(\psi))$ owing to the dominated convergence theorem.
(i) We have for all $\omega\in\Omega$,
$|\hat{\vartheta}_{N}^{*}(\omega)-\vartheta^{*}|\leq\sup_{u\in\mathcal{D}(\psi)}\,|F(Bu)+\psi(u)-\big{(}\hat{F}_{N}(Bu,\omega)-\psi(u)\big{)}|\leq\sup_{w\in B(\mathcal{D}(\psi))}|F(w)-\hat{F}_{N}(w,\omega)|.$
Combined with (3.2), we find that $\hat{\vartheta}_{N}^{*}(\omega)\to\vartheta^{*}$ as $N\to\infty$ for all $\omega\in\Omega\setminus\Omega_{0}$.
(ii) Fix $\omega\in\Omega\setminus\Omega_{0}$. Since $(u_{N}^{*}(\omega))$ is bounded, it has a weak accumulation point. Let $\bar{u}(\omega)\in U$ be a weak accumulation point of $(u_{N}^{*}(\omega))$. By assumption, there exists a subsequence $(u_{N}^{*}(\omega))_{\mathcal{N}(\omega)}$ of $(u_{N}^{*}(\omega))$ such that $u_{N}^{*}(\omega)\rightharpoonup\bar{u}(\omega)\in U$ as $\mathcal{N}(\omega)\ni N\to\infty$. Fix $u\in\mathcal{D}(\psi)$. We have
$\hat{\vartheta}_{N}^{*}(\omega)=\hat{F}_{N}(Bu_{N}^{*}(\omega),\omega)+\psi(u_{N}^{*}(\omega))\leq\hat{F}_{N}(Bu,\omega)+\psi(u).$
Using the convergence statement in (3.2), and $Bu_{N}^{*}(\omega)\to B\bar{u}(\omega)$ as $\mathcal{N}(\omega)\ni N\to\infty$, we obtain $\hat{F}_{N}(Bu_{N}^{*}(\omega))\to F(B\bar{u}(\omega))$ and $\hat{F}_{N}(Bu,\omega)\to F(Bu)$ as $\mathcal{N}(\omega)\ni N\to\infty$. Combined with the fact that $\psi$ is weakly lower semicontinuous,
$F(B\bar{u}(\omega))+\psi(\bar{u}(\omega))$ $\leq\liminf_{\mathcal{N}(\omega)\ni N\to\infty}\,\{\,\hat{F}_{N}(Bu_{N}^{*}(\omega))+\psi(u_{N}^{*}(\omega))\,\}\leq\lim_{\mathcal{N}(\omega)\ni N\to\infty}\,\hat{F}_{N}(Bu)+\psi(u)$
$=F(Bu)+\psi(u).$
Since $\Omega\setminus\Omega_{0}\in\mathcal{F}$ occurs w.p. $1$, we obtain the assertion. ∎
3.3 Consistency of SAA critical points
We establish the asymptotic consistency of the SAA gap functional evaluated as approximate SAA critical points as well as that of SAA critical points.
Theorem 3.4.
Let Assumptions 1 and 2 hold, and let $(\varepsilon_{N})\subset[0,\infty)$ be a sequence with $\varepsilon_{N}\to 0$ as $N\to\infty$. Let $(\bar{u}_{N})\subset\mathcal{D}(\psi)$ be a sequence of random vectors with $\hat{\Psi}_{N}(\bar{u}_{N})\leq\varepsilon_{N}$ w.p. $1$. Then (i) $\Psi(\bar{u}_{N})\to 0$ as $N\to\infty$ w.p. $1$. (ii) For almost all $\omega\in\Omega$, $(\bar{u}_{N}(\omega))$ has at least one weak accumulation point in $U$ and every such point is a zero of $\Psi$.
Proof.
Before we establish parts (i) and (ii), let us note that for all $u\in\mathcal{D}(\psi)$,
$\Psi(u)$ $\leq\hat{\Psi}_{N}(u)+\sup_{v\in\mathcal{D}(\psi)}\,(B^{*}\mathrm{D}F(Bu)-B^{*}\mathrm{D}\hat{F}_{N}(Bu),u-v)_{U}.$ (3.3)
Hence, we have for all $u\in\mathcal{D}(\psi)$,
[…]
Therefore, the uniform law of large numbers [14, Cor. 4:1] ensures the existence of a null set $\Omega_{0}\subset\Omega$ with $\Omega_{0}\in\mathcal{F}$ such that for all $\omega\in\Omega\setminus\Omega_{0}$, we have
$\sup_{(v,w)\in B(\mathcal{D}(\psi))\times B(\mathcal{D}(\psi))}\,\langle\mathrm{D}F(w)-\mathrm{D}\hat{F}_{N}(w,\omega),w-v\rangle_{{W}^{*},W}\to 0\quad\text{as}\quad N\to\infty.$
(i) Using $\hat{\Psi}_{N}(\bar{u}_{N})\leq\varepsilon_{N}$, the error bound (3.4), and the above convergence statement, we obtain $\Psi(\bar{u}_{N})\to 0$ as $N\to\infty$ w.p. $1$.
(ii) Fix $\omega\in\Omega\setminus\Omega_{0}$ such that $\Psi(\bar{u}_{N}(\omega))\to 0$ as $N\to\infty$. Since $(\bar{u}_{N}(\omega))$ is bounded, it has a weak accumulation point. If $\bar{u}_{N}(\omega)\rightharpoonup\bar{u}(\omega)$ as $\mathcal{N}(\omega)\ni N\to\infty$, then the weak lower semicontinuity of $\Psi$ established in Proposition 3.2) and $\Psi(\bar{u}_{N}(\omega))\to 0$ as $N\to\infty$ imply the assertion. ∎
3.4 Sample size estimates for SAA critical points
We establish expectation bounds and sample size estimates. Our analysis is inspired by that of [34].
For this purpose, we formulate a Lipschitz continuity assumption and a light-tailed condition on the integrand’s gradient, which are a typical assumptions in related contexts.
Assumption 3 (Lipschitz-type condition on integrand’s derivative, sub-Gaussian derivatives).
  1. (a)
    For an integrable random variable $L_{\mathrm{D}f}:\Xi\to[0,\infty)$,
    $\|\nabla g_{\xi}(u_{2})-\nabla g_{\xi}(u_{1})\|_{U}\leq L_{\mathrm{D}f}(\xi)\|Bu_{2}-Bu_{1}\|_{W}\quad\text{for all}\quad u_{1},u_{2}\in\mathcal{D}(\psi),\quad\xi\in\Xi.$
  2. (b)
    For some constant $\tau_{\mathrm{D}f}>0$,
    $\mathbb{E}[\exp(\tau_{\mathrm{D}f}^{-2}\|\nabla g_{\boldsymbol{\xi}}(u)-\mathbb{E}[\nabla g_{\boldsymbol{\xi}}(u)]\|_{U}^{2})]\leq\mathrm{e}\quad\text{for all}\quad u\in\mathcal{D}(\psi).$
3 (b) is, for example, satisfied if there exists a constant $C>0$ such that $\|\nabla g_{\xi}(u)\|_{U}\leq C$ for all $(u,\xi)\in\mathcal{D}(\psi)\times\Xi$. We denote by $r_{\text{ad}}$ the diameter of $\mathcal{D}(\psi)$.
Theorem 3.5.
Let Assumptions 1, 2 and 3 hold. For each $N\in\mathbb{N}$, let $u_{N}\colon\Omega\to U$ be measurable. Then, for all $N\in\mathbb{N}$ and $\nu>0$,
$\mathbb{E}[\Psi(u_{N})]\leq\mathbb{E}[\hat{\Psi}_{N}(u_{N})]+2L_{\mathrm{D}f}r_{\text{ad}}\nu+\tfrac{\sqrt{3}\tau_{\mathrm{D}f}r_{\text{ad}}}{\sqrt{N}}\big{(}\mathcal{N}(\nu;B(\mathcal{D}(\psi)))\big{)}^{1/2}.$ (3.5)
Proof.
Using (3.3), we have
$\mathbb{E}[\Psi(u_{N})]\leq\mathbb{E}[\hat{\Psi}_{N}(u_{N})]+r_{\text{ad}}\mathbb{E}[\sup_{v\in B(\mathcal{D}(\psi))}\,\|B^{*}\mathrm{D}F(v)-B^{*}\mathrm{D}\hat{F}_{N}(v)\|_{U}].$
Applying Proposition A.1 to $\mathsf{G}(v,\xi)\coloneqq B^{*}\mathrm{D}_{w}f(v,\xi)$, we obtain
$\mathbb{E}[\sup_{v\in B(\mathcal{D}(\psi))}\|B^{*}\mathrm{D}F(v)-B^{*}\mathrm{D}\hat{F}_{N}(v)\|_{U}]\leq 2L_{\mathrm{D}f}\nu+\tfrac{\sqrt{3}\tau_{\mathrm{D}f}}{\sqrt{N}}\big{(}\mathcal{N}(\nu;B(\mathcal{D}(\psi)))\big{)}^{1/2}.$
Using Theorem 3.5, we establish sample size estimates.
Corollary 3.6.
Let the hypotheses of Theorem 3.5 hold. If $\varepsilon>0$ and
$N\geq\frac{12r_{\text{ad}}^{2}\tau_{\mathrm{D}f}^{2}}{\varepsilon^{2}}\cdot\mathcal{N}(\varepsilon/(4r_{\text{ad}}L_{\mathrm{D}f});B(\mathcal{D}(\psi))),$
then $\mathbb{E}[\Psi(u_{N})]\leq\mathbb{E}[\hat{\Psi}_{N}(u_{N})]+\varepsilon$.
Proof.
We use (3.5) with $\nu=\varepsilon/(4r_{\text{ad}}L_{\mathrm{D}f})$. We have $22L_{\mathrm{D}f}\nu=\varepsilon/$ 2. The choice of $N$ ensures
$\tfrac{\sqrt{3}\tau_{\mathrm{D}f}r_{\text{ad}}}{\sqrt{N}}\big{(}\mathcal{N}(\nu;B(\mathcal{D}(\psi)))\big{)}^{1/2}\leq\varepsilon/2.$
The sample size estimates in Corollary 3.6 are based on the covering numbers of $B(\mathcal{D}(\psi))$. A typical example in PDE-constrained optimization models $B$ as the adjoint operator of the embedding $H^{1}(D)\hookrightarrow L^{2}(D)$. Combining the covering numbers for Sobolev function classes in [6, Thm. 1.7] and duality of metric entropy [3, p. 1315], covering numbers of $B(\mathcal{D}(\psi))$ can be derived.
3.5 Expectation bounds for convex problems
For convex problems, we demonstrate nonasymptotic expectation bounds for the distance between SAA solutions and the true solution, the gap functional, and the optimality gap. These statements hold true without the compactness of the linear bounded operator $B$. In the following, let $u^{*}$ be a solution to ($\mathcal{P}$) and for each $N\in\mathbb{N}$, let $u_{N}^{*}$ be a measurable solution to ($\mathcal{P}_{N}$).
Assumption 4 (Hölder-type inequality, gradient regularity, growth, and standard deviation).
  1. (a)
    The space $\mathcal{U}$ is a Banach space with $\mathcal{U}^{*}\hookrightarrow U\hookrightarrow\mathcal{U}$, $\mathcal{H}$ is a separable Hilbert space with $\mathcal{H}\hookrightarrow\mathcal{U}^{*}$, and
    $|(v,u)_{U}|\leq\|v\|_{\mathcal{U}^{*}}\|u\|_{\mathcal{U}}\quad\text{for all}\quad v\in\mathcal{U}^{*},\quad u\in U.$
  2. (b)
    For two constants $\theta\in(0,\infty)$ and $\varrho\in[2,\infty)$, it holds that
    $(B^{*}\mathrm{D}F(Bu^{*}),u-u^{*})_{U}+\psi(u)-\psi(u^{*})\geq\theta\|u-u^{*}\|_{\mathcal{U}}^{\varrho}\quad\text{for all}\quad u\in U.$
  3. (c)
    The mapping $\Xi\ni\xi\mapsto\nabla g_{\xi}(u^{*})\in\mathcal{H}$ is integrable, and standard deviation-type constant
    $\sigma_{\mathrm{D}f}\coloneqq(\mathbb{E}[\|\nabla g_{\boldsymbol{\xi}}(u^{*})-\mathbb{E}[\nabla g_{\boldsymbol{\xi}}(u^{*})]\|_{\mathcal{H}}^{2}])^{1/2}$
    is finite.
4 (b) and its variants have been utilized, for example, in [13] and, in particular, ensures that $u^{*}$ is the unique solution to ($\mathcal{P}$).
We state the section’s main result. If 4 (a) holds true, let $C_{\mathcal{H};\mathcal{U}^{*}}\in(0,\infty)$ be the embedding constant of the embedding $\mathcal{H}\hookrightarrow\mathcal{U}^{*}$. If $\varrho=2$ in Assumption 4, the following result ensures the typical Monte Carlo convergence rate, $1/\sqrt{N}$, for the expected error $\mathbb{E}[\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}]$.
Theorem 3.7.
If Assumptions 1, 2 and 4 hold, and $f(\cdot,\xi)$ is convex for all $\xi\in\Xi$, then for all $N\in\mathbb{N}$,
$\theta^{2}\mathbb{E}[\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}^{2(\varrho-1)}]\leq(1/N)C_{\mathcal{H};\mathcal{U}^{*}}^{2}\sigma_{\mathrm{D}f}^{2}.$
Proof.
The proof is inspired by those of Lemma 6 and Theorem 3 in [21]. Using $-B^{*}\mathrm{D}F(Bu^{*})\in\partial\psi(u^{*})$ and 4 (b), we obtain
$(B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}\hat{F}_{N}(Bu_{N}^{*}),u_{N}^{*}-u^{*})_{U}\geq\theta\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}^{\varrho}.$
Since $u\mapsto\hat{F}_{N}(Bu)$ is continuously differentiable on $U_{0}$ (see Proposition 3.1), and convex, we have
$(B^{*}\mathrm{D}\hat{F}_{N}(Bu_{N}^{*})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*}),u_{N}^{*}-u^{*})_{U}\geq 0.$
Adding both estimates ensures
$(B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*}),u_{N}^{*}-u^{*})_{U}\geq\theta\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}^{\varrho}.$
Hence
$\|B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*})\|_{\mathcal{U}^{*}}\geq\theta\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}^{\varrho-1}.$
Taking squares and using the continuity of the embedding $\mathcal{H}\hookrightarrow\mathcal{U}^{*}$, we obtain
$C_{\mathcal{H};\mathcal{U}^{*}}^{2}\|B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*})\|_{\mathcal{H}}^{2}\geq\theta^{2}\|u_{N}^{*}-u^{*}\|_{\mathcal{U}}^{2(\varrho-1)}.$
Since $\mathbb{E}[B^{*}\mathrm{D}_{w}f(Bu^{*},\boldsymbol{\xi})]=B^{*}\mathrm{D}F(Bu^{*})$ (see Proposition 3.1), $B^{*}\mathrm{D}_{w}f(Bu^{*},\boldsymbol{\xi}^{i})$, $i=1,2,\ldots$ are independent and identically distributed, and $\mathcal{H}$ is a separable Hilbert space, we obtain
$\mathbb{E}[\|B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*})\|_{\mathcal{H}}^{2}]=(1/N)\sigma_{\mathrm{D}f}^{2}.$ (3.6)
This implies the expectation bound. ∎
Using this result, we further deduce superconvergence properties of the objective function values and the gap functional evaluated at SAA solutions. We recall that $G$ denotes the objective function of ($\mathcal{P}$).
Corollary 3.8.
Let the hypotheses of Theorem 3.7 hold with $\varrho=2$, and assume that there exists $L_{\mathrm{D}f,\mathcal{U}}>0$ such that
$\|\nabla g_{\xi}(u_{2})-\nabla g_{\xi}(u_{1})\|_{\mathcal{U}^{*}}\leq L_{\mathrm{D}f,\mathcal{U}}\|u_{2}-u_{1}\|_{\mathcal{U}}\quad\text{for all}\quad u_{1},u_{2}\in\mathcal{D}(\psi),\quad\xi\in\Xi.$ (3.7)
Then, for all $N\in\mathbb{N}$, we have
$\mathbb{E}[G(u_{N}^{*})-G(u^{*})]\leq\mathbb{E}[\Psi(u^{*}_{N})]\leq\bigg{(}1+\bigg{(}\frac{1}{\theta}+\frac{L_{\mathrm{D}f,\mathcal{U}}}{\theta^{2}}\bigg{)}^{2}+\frac{4L_{\mathrm{D}f,\mathcal{U}}^{2}}{\theta^{2}}\bigg{)}\frac{C_{\mathcal{H};\mathcal{U}^{*}}^{2}\sigma^{2}_{\mathrm{D}f}}{N}.$
Proof.
Since $G(u)-G(u^{*})\leq\Psi(u)$ for all $u\in\mathcal{D}(\psi)$ (see, for example, [13, Thm. 2.4]), it suffices to show the claimed estimate for the gap functional. For this purpose, note that there exists $v_{N}\in\mathcal{D}(\psi)$ such that
$\Psi(u^{*}_{N})$ $=(B^{*}\mathrm{D}F(Bu^{*}_{N}),u^{*}_{N}-v_{N})_{U}+\psi(u^{*}_{N})-\psi(v_{N})\leq(B^{*}\mathrm{D}F(Bu^{*}_{N})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*}_{N}),u^{*}_{N}-v_{N})_{U}$
$\leq(1/2)\|B^{*}\mathrm{D}F(Bu^{*}_{N})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*}_{N})\|_{\mathcal{U}^{*}}^{2}+(1/2)\|u^{*}_{N}-v_{N}\|_{\mathcal{U}}^{2},$
where the first inequality follows from the optimality of $u_{N}^{*}$. Using (3.7), we have
$\|B^{*}\mathrm{D}F(Bu^{*}_{N})-B^{*}\mathrm{D}\hat{F}_{N}(Bu^{*}_{N})\|_{\mathcal{U}^{*}}\leq 2L_{\mathrm{D}f,\mathcal{U}}\|u^{*}_{N}-u^{*}\|_{\mathcal{U}}+C_{\mathcal{H};\mathcal{U}^{*}}\|B^{*}\mathrm{D}F(Bu^{*})-\mathrm{D}\hat{F}_{N}(Bu^{*})\|_{\mathcal{H}}.$
Following the proof of Lemma 4.8 in [13], we arrive at
$\theta\|v_{N}-u^{*}\|_{\mathcal{U}}^{2}$ $\leq(B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}F(Bu^{*}_{N}),v_{N}-u^{*})_{U}\leq\|B^{*}\mathrm{D}F(Bu^{*})-B^{*}\mathrm{D}F(Bu^{*}_{N})\|_{\mathcal{U}^{*}}\|v_{N}-u^{*}\|_{\mathcal{U}}$
$\leq L_{\mathrm{D}f,\mathcal{U}}\|u^{*}_{N}-u^{*}\|_{\mathcal{U}}\|v_{N}-u^{*}\|_{\mathcal{U}},$
where the final inequality is due to (3.7). Putting together the pieces, we have
$\Psi(u^{*}_{N})\leq\frac{1}{2}\bigg{(}1+\frac{L_{\mathrm{D}f,\mathcal{U}}}{\theta}\bigg{)}^{2}\|u^{*}_{N}-u^{*}\|_{\mathcal{U}}^{2}+4L_{\mathrm{D}f,\mathcal{U}}^{2}\|u^{*}_{N}-u^{*}\|_{\mathcal{U}}^{2}+C_{\mathcal{H};\mathcal{U}^{*}}^{2}\|B^{*}\mathrm{D}F(Bu^{*})-\mathrm{D}\hat{F}_{N}(Bu^{*})\|_{\mathcal{H}}^{2}.$
Taking expectations, and using (3.6) and Theorem 3.7, we obtain the expectation bound. ∎
4 Application to bang-bang optimal control under uncertainty
In this section, we apply our results to risk-neutral, nonsmooth PDE-constrained minimization problems of the form
$\min_{u\in L^{2}(D)}\,\mathbb{E}[J(\imath y(\boldsymbol{\xi}))]+\beta\|u\|_{L^{1}(D)}\,\quad\text{s.t.}\quad u\in[\mathfrak{l},\mathfrak{u}],$ (4.1)
where $D\subset\mathbb{R}^{d}$$d\in\{1,2,3\}$, is a bounded, polyhedral domain, $\beta\geq 0$, $J$ is a smooth fidelity term defined on $L^{2}(D)$ and $\imath\colon H^{1}_{0}(D)\to L^{2}(D)$ denotes the compact embedding. Moreover,
$[\mathfrak{l},\mathfrak{u}]\coloneqq\{\,v\in L^{2}(D)\;|\;\mathfrak{l}(x)\leq v(x)\leq\mathfrak{u}(x)\text{ a.e. }x\in D\,\},\;\;\text{where}\;\;\mathfrak{l},\mathfrak{u}\in L^{\infty}(D),\;\;\mathfrak{l}(x)\leq 0\leq\mathfrak{u}(x)\text{ a.e. }x\in D.$
The deterministic control $u\in[\mathfrak{l},\mathfrak{u}]$ is coupled with the random state variable $y(\xi)\in H^{1}_{0}(D)$ via a parametrized elliptic PDE
$(\kappa(\xi)\nabla y,\nabla v)_{L^{2}(D)^{d}}+(E(u,y),v)_{L^{2}(D)^{d}}=(b(\xi),v)_{L^{2}(D)}\quad\text{for all}\quad v\in H_{0}^{1}(D),$ (4.2)
which the state satisfies for every $\xi\in\Xi$. Throughout this section, we consider
$U=L^{2}(D),\quad W=H^{-1}(D),\quad\mathcal{H}=H^{2}(D),\quad\mathcal{U}=L^{1}(D),\quad\text{and}\quad B=\imath^{*}.$
In the following, we show that both linear, that is, $E(u,y)=-u$, as well as bilinear, that is, $E(u,y)=uy$, control problems fit into the theoretical framework ($\mathcal{P}$) defining
$\psi(u)=\beta\|u\|_{L^{1}(D)}+I_{[\mathfrak{l},\mathfrak{u}]}(u)$
as well as introducing the integrand
$f(w,\xi)\coloneqq J(B^{*}S(w,\xi)),\quad\text{where}\quad S\colon B(U_{0})\times\Xi\to H^{1}_{0}(D)$ (4.3)
denotes a suitable parametrized control-to-state operator defined on the image of a $L^{2}(D)$-neighborhood $U_{0}$ of $\mathcal{D}(\psi)=[\mathfrak{l},\mathfrak{u}]$. For both problem classes, the following assumptions are made.
Assumption 5 (Fidelity term, PDE right-hand side, random diffusion coefficient).
  1. (a)
    The function $J:L^{2}(D)\to[0,\infty)$ is convex and continuously differentiable. Its gradient $\nabla J$ is Lipschitz continuous with Lipschitz constant $\ell\in(0,\infty)$. Moreover $J(v)\leq\varrho(\|v\|_{L^{2}(D)})$ for all $v\in L^{2}(D)$, where $\varrho:[0,\infty)\to[0,\infty)$ is a polynomial and nondecreasing.
  2. (b)
    The map $b:\Xi\to L^{2}(D)$ is measurable and there exists $b_{\sup}^{*}>0$ such that $\|b(\xi)\|_{L^{2}(D)}\leq b_{\sup}^{*}$ for all $\xi\in\Xi$.
  3. (c)
    The mapping $\kappa:\Xi\to C^{1}(\bar{D})$ is measurable and $\mathbb{E}[\|\kappa(\boldsymbol{\xi})\|_{C^{1}(\bar{D})}^{p}]<\infty$ for all $p\in[1,\infty)$. Moreover, there exists $\kappa_{\inf}^{*}>0$ such that $\kappa(\xi)(x)\geq\kappa_{\inf}^{*}$ for all $x\in D$ and $\xi\in\Xi$.
We point out that 5 (a) implies
$\|\nabla J(v)\|_{L^{2}(D)}^{2}\leq 2\ell J(v)\leq 2\ell\varrho(\|v\|_{L^{2}(D})\ \quad\text{for all}\quad v\in L^{2}(D).$ (4.4)
Moreover, 1 (a), 1 (b) and 1 (c) are clearly satisfied. For abbreviation, let $C_{D}$ be the Friedrichs constant of the domain $D$, which equals the operator norm of $\imath$.
4.1 Risk-neutral affine-linear PDE-constrained optimization
We first consider (4.1) governed by a class of affine-linear, parameterized PDEs, that is, $E(u,y)=-u$. For simplicity, set $b=0$. We consider the operator
$A\colon C^{1}(\bar{D})\times H^{-1}(D)\times H^{1}_{0}(D)\to H^{-1}(D),$
where
$\langle A(\kappa,w,y),v\rangle_{H^{-1}(D),H_{0}^{1}(D)}\coloneqq(\kappa\nabla y,\nabla v)_{L^{2}(D)^{d}}-\langle w,v\rangle_{H^{-1}(D),H_{0}^{1}(D)}\quad\text{for all}\quad v\in H_{0}^{1}(D),$
and the associated parametrized equation
$\text{find}~{}y\in H^{1}_{0}(D)\colon\quad A(\kappa,w,u)=0,\quad\text{where}\quad(\kappa,w)\in\mathfrak{P}_{\text{ad}}\coloneqq\left\{\,\kappa\in C^{1}(\bar{D})\;|\;\kappa>\kappa^{*}_{\inf}/2\,\right\}\times H^{-1}(D).$ (4.5)
The following lemma is a direct consequence of the Lax–Milgram lemma, the implicit function theorem as well as standard regularity results for elliptic PDEs. Its proof is omitted for the sake of brevity.
Lemma 4.1.
Let Assumption 5 hold with $b=0$. For every $(\kappa,w)\in\mathfrak{P}_{\text{ad}}$, there exists a unique $y=y(\kappa,w)\in H^{1}_{0}(D)$ satisfying (4.5). The mapping $\mathfrak{P}_{\text{ad}}\ni(\kappa,w)\mapsto y(\kappa,w)$ is infinitely many times continuously differentiable, and we have
$\|y(\kappa,w)\|_{H_{0}^{1}(D)}\leq(2/\kappa_{\inf}^{*})\|w\|_{H^{-1}(D)}\quad\text{for all}\quad(\kappa,w)\in\mathfrak{P}_{\text{ad}}.$ (4.6)
Moreover, if $u\in L^{2}(D)$, then $y(\kappa,Bu)\in H^{1}_{0}(D)\cap H^{2}(D)$, and there exists a constant $C_{H_{2}}>0$ such that
$\|y(\kappa,Bu)\|_{H^{2}(D)}\leq\frac{C_{H^{2}}\|\kappa\|_{C^{1}(\bar{D})}^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\|u\|_{L^{2}(D)}\quad\text{for all}\quad(\kappa,Bu)\in\mathfrak{P}_{\text{ad}},\quad u\in L^{2}(D).$ (4.7)
Finally, the mapping
$\left\{\,\kappa\in C^{1}(\bar{D})\;|\;\kappa>\kappa^{*}_{\inf}/2\,\right\}\times L^{2}(D)\times\ni(\kappa,u)\mapsto y(\kappa,b,u)\in H^{1}_{0}(D)\cap H^{2}(D)$
is continuous.
In the following, let $U_{0}\subset L^{2}(D)$ be an arbitrary but fixed convex, bounded neighborhood of $\mathcal{D}(\psi)$, and let $C_{U_{0}}>0$ be such that $\|u\|_{L^{2}(D)}\leq C_{U_{0}}$ for all $u\in U_{0}$. Defining the parametrized control-to-state operator
$S\colon B(U_{0})\times\Xi\to H^{1}_{0}(D),\quad\text{where}\quad S(w,\xi)\coloneqq y(\kappa(\xi),w),$
we can verify the remaining assumptions in the linear case.
Lemma 4.2.
If Assumption 5 hold with $b=0$, then $f\colon B(U_{0})\times\Xi\to\mathbb{R}$ in (4.3) is Carathéodory, and
$0\leq f(Bu,\xi)\leq\varrho(C_{U_{0}}C^{2}_{D}(2/\kappa^{*}_{\inf}))\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi.$
Proof.
The Carathéodory property of $f$ follows from the continuity of $J$ and that of the mapping $y$ from Lemma 4.1. Moreover, leveraging 5 (a) and (4.6) yields
$0\leq f(Bu,\xi)\leq\varrho(C^{2}_{D}(2/\kappa^{*}_{\inf})\|u\|_{L^{2}(D)})\leq\varrho(C_{U_{0}}C^{2}_{D}(2/\kappa^{*}_{\inf}))\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi.$
Next, we verify Assumptions 2 and 5 (a).
Lemma 4.3.
Let Assumption 5 hold with $b=0$. We define
$\mathrm{D}_{w}f\colon B(U_{0})\times\Xi\to H^{1}_{0}(D),\quad\text{where}\quad\mathrm{D}_{w}f(w,\xi)\coloneqq S\left(B\nabla J(B^{*}S(w,\xi)),\xi\right).$
The mapping $\mathrm{D}_{w}f$ is Carathédoroy and, for every $\xi\in\Xi$, the function
$g_{\xi}\colon L^{2}(D)\to\mathbb{R},\quad\text{where}\quad g_{\xi}(u)\coloneqq J(B^{*}S(Bu,\xi))$
is continuously differentiable on $L^{2}(D)$ with $\nabla g_{\xi}(u)=B^{*}\mathrm{D}_{w}f(Bu,\xi)$. Moreover, we have the estimates
$\|\nabla g_{\xi}(u)\|_{L^{2}(D)}\leq C_{D}\|\mathrm{D}_{w}f(Bu,\xi)\|_{H^{1}_{0}(D)}\leq C^{2}_{D}(2/\kappa_{\inf}^{*})\sqrt{2\ell\varrho(C^{2}_{D}(2/\kappa_{\inf}^{*})C_{U_{0}})}\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi.$
Finally, we have
$\|\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})\|_{L^{2}(D)}\leq C^{4}_{D}\ell(2/\kappa_{\inf}^{*})^{2}\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}\quad\text{for all}\quad u_{1},u_{2}\in\mathcal{D}(\psi),\quad\xi\in\Xi.$
Proof.
The Carathéodory property of $\mathrm{D}_{w}f$ follows analogously to Lemma 4.2. Moreover, the differentiability of $g_{\xi}$ and the representation of its gradient are direct consequences of the chain rule as well as standard adjoint calculus, respectively. Using this representation and (4.6), we obtain
$\|\nabla g_{\xi}(u)\|_{L^{2}(D)}\leq C_{D}\|\mathrm{D}_{w}f(Bu,\xi)\|_{H^{1}_{0}(D)}\leq C^{2}_{D}(2/\kappa_{\inf}^{*})\|\nabla J(B^{*}S(Bu,\xi))\|_{L^{2}(D)}.$
Furthermore, by applying (4.4), we have
$\|\nabla J(B^{*}S(Bu,\xi))\|_{L^{2}(D)}^{2}\leq 2\ell\varrho(\|B^{*}S(Bu,\xi)\|_{L^{2}(D)})\leq 2\ell\varrho(C^{2}_{D}(2/\kappa_{\inf}^{*})C_{U_{0}}),$ (4.8)
where the last inequality follows again by using (4.6). Combining both inequalities yields the desired estimate. The Lipschitz estimate for $\nabla g_{\xi}$ follows analogously noting that
$\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})=B^{*}S\big{(}B\nabla J(B^{*}S(Bu_{1},\xi))-B^{*}\nabla J(B^{*}S(Bu_{2},\xi)),\xi\big{)},$
and thus
$\|\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})\|_{L^{2}(D)}$ $\leq C^{2}_{D}(2/\kappa_{\inf}^{*})\|\nabla J(B^{*}S(Bu_{1},\xi))-\nabla J(B^{*}S(Bu_{2},\xi))\|_{L^{2}(D)}$
$\leq C^{3}_{D}\ell(2/\kappa_{\inf}^{*})\|S(Bu_{1},\xi)-S(Bu_{2},\xi)\|_{H^{1}_{0}(D)}$
$\leq C^{4}_{D}\ell(2/\kappa_{\inf}^{*})^{2}\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}$
for all $u_{1},u_{2}\in\mathcal{D}(\psi)$ and $\xi\in\Xi$. ∎
Next, we discuss sufficient conditions for Assumption 4 with the particular choice of $\mathcal{U}=L^{1}(D)$ and $\mathcal{H}=H^{2}(D)$. In this case, 4 (a) is satisfied. In the following, we denote by $u^{*}$ a solution to (4.1). We recall the identity $\mathrm{D}F(Bu^{*})=\mathbb{E}[\mathrm{D}_{w}f(Bu^{*},\boldsymbol{\xi})]$.
Proposition 4.4.
Let Assumption 5 hold with $b=0$. Then, the mapping $\Xi\ni\xi\mapsto\mathrm{D}_{w}f(Bu^{*},\xi)\in H^{2}(D)$ is measurable, and
$\mathbb{E}[\|\mathrm{D}_{w}f(Bu^{*},\boldsymbol{\xi})-\mathbb{E}[\mathrm{D}_{w}f(Bu^{*},\boldsymbol{\xi})]\|_{H^{2}(D)}^{2}]\leq\frac{2\ell C^{2}_{H^{2}}\mathbb{E}[\|\kappa(\boldsymbol{\xi})\|_{C^{1}(\bar{D})}^{6}]}{(\kappa^{*}_{\inf}/2)^{8}}\varrho(C^{2}_{D}(2/\kappa_{\inf}^{*})C_{U_{0}})<\infty.$
If, moreover, there exist constants  $C>0$ and $\nu\in(0,1]$ such that
$\operatorname{meas}\big{\{}\,x\in D\;\colon\;\big{|}|[\mathrm{D}F(Bu^{*})](x)|-\beta\big{|}\leq\varepsilon\,\big{\}}\leq C\varepsilon^{\nu}\quad\text{for all}\quad\varepsilon>0,$
then there exists $\theta>0$ such that
$(B^{*}\mathrm{D}F(Bu^{*}),u-u^{*})_{L^{2}(D)}+\psi(u)-\psi(u^{*})\geq\theta\|u-u^{*}\|_{\mathcal{U}}^{1+1/\nu}\quad\text{for all}\quad u\in L^{2}(D).$
Proof.
Lemma 4.1 ensures $\mathrm{D}_{w}f(Bu^{*},\xi)\in H^{2}(D)$ as well as the claimed measurability. Using (4.7) and (4.8), we arrive at
$\|\mathrm{D}_{w}f(Bu^{*},\xi)\|_{H^{2}(D)}^{2}$ $\leq\frac{C^{2}_{H^{2}}\|\kappa(\xi)\|_{C^{1}(\bar{D})}^{6}}{(\kappa^{*}_{\inf}/2)^{8}}\|\nabla J(B^{*}S(Bu^{*},\xi))\|_{L^{2}(D)}^{2}\leq\frac{2\ell C^{2}_{H^{2}}\|\kappa(\xi)\|_{C^{1}(\bar{D})}^{6}}{(\kappa^{*}_{\inf}/2)^{8}}\varrho(C^{2}_{D}(2/\kappa_{\inf}^{*})C_{U_{0}}).$
Taking expectations yields the expectation bound. Finally, the growth condition follows analogously to Proposition 5.5 in [13]. ∎
Finally, we comment on the uniform Lipschitz continuity assumed in Corollary 3.8.
Lemma 4.5.
Let Assumption 5 hold with $b=0$. Assume that the mapping $\kappa:\Xi\to C^{1}(\bar{D})$ from 5 (c) satisfies $\|\kappa(\xi)\|_{C^{1}(\bar{D})}\leq\kappa_{\sup}^{*}$ all $\xi\in\Xi$ and some $\kappa_{\sup}^{*}>0$. Then, there holds
$\|\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})\|_{L^{\infty}(D)}\leq\ell\bigg{(}C_{H^{2};L^{\infty}}\frac{C_{H^{2}}(\kappa_{\sup}^{*})^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\bigg{)}^{2}\|u_{1}-u_{2}\|_{L^{1}(D)}\quad\text{for all}\quad u_{1},u_{2}\in\mathcal{D}(\psi),\quad\xi\in\Xi$
where $C_{H^{2}}>0$ is the constant from Lemma 4.1, and $C_{H^{2};L^{\infty}}>0$ denotes the embedding constant of $H^{2}(D)\hookrightarrow L^{\infty}(D)$.
Proof.
Similarly to the proof of Lemma 4.3, we obtain
$\|\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})\|_{L^{\infty}(D)}$ $\leq C_{H^{2};L^{\infty}}\frac{C_{H^{2}}(\kappa_{\sup}^{*})^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\|\nabla J(B^{*}S(Bu_{1},\xi))-\nabla J(B^{*}S(Bu_{2},\xi))\|_{L^{2}(D)}$
$\leq\ell C_{H^{2};L^{\infty}}\frac{C_{H^{2}}(\kappa_{\sup}^{*})^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\|S(Bu_{1},\xi)-S(Bu_{2},\xi)\|_{L^{2}(D)}.$
using (4.7) as well as the assumed boundedness of $\kappa$. Furthermore, by standard adjoint arguments and by again exploiting (4.7), we arrive at
$\|S(Bu_{1},\xi)-S(Bu_{2},\xi)\|_{L^{2}(D)}$ $=\sup_{\|v\|_{L^{2}(D)}\leq 1}(S(Bu_{1}-Bu_{2},\xi),v)_{L^{2}(D)}=\sup_{\|v\|_{L^{2}(D)}\leq 1}(S(Bv,\xi),u_{1}-u_{2})_{L^{2}(D)}$
$\leq C_{H^{2};L^{\infty}}\sup_{\|v\|_{L^{2}(D)}\leq 1}\|S(Bv,\xi)\|_{H^{2}(D)}\|u_{1}-u_{2}\|_{L^{1}(D)}$
$\leq C_{H^{2};L^{\infty}}\frac{C_{H^{2}}(\kappa_{\sup}^{*})^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\|u_{1}-u_{2}\|_{L^{1}(D)}.$
Combining both observations yields the desired estimate. ∎
4.2 Risk-neutral bilinear PDE-constrained optimization
Next, we show that bilinear control problems, that is, $E(u,y)=uy$, also fit our abstract setting. For this purpose, we assume $\mathfrak{l}(x)\geq 0$ for a.e. $x\in D$. Similar to the previous example, consider the operator
$A\colon C^{1}(\bar{D})\times L^{2}(D)\times H^{1}_{0}(D)\times L^{2}(D)\to H^{-1}(D),$
where
$\langle A(\kappa,b,y,u),v\rangle_{H^{-1}(D),H_{0}^{1}(D)}\coloneqq(\kappa\nabla y,\nabla v)_{L^{2}(D)^{d}}+(yu,v)_{L^{2}(D)}-(b,v)_{L^{2}(D)}\quad\text{for all}\quad v\in H_{0}^{1}(D).$
We readily verify that $A$ is infinitely many times continuously differentiable. Now, define the open set
$\mathfrak{P}_{\text{ad}}\coloneqq\left\{\,\kappa\in C^{1}(\bar{D})\;|\;\kappa>\kappa^{*}_{\inf}/2\,\right\}\times L^{2}(D)\times V_{\text{ad}},$
where
$V_{\text{ad}}\coloneqq\mathcal{D}(\psi)+(\kappa_{\inf}^{*}/(4C_{H_{0}^{1};L^{4}}^{2}))\cdot\mathbb{B}_{L^{2}(D)}(0).$
Moreover, $\mathbb{B}_{L^{2}(D)}(0)$ is the open $L^{2}(D)$-unit ball, and $C_{H_{0}^{1};L^{4}}>0$ is the embedding constant of $H^{1}_{0}(D)\hookrightarrow L^{4}(D)$. We consider the equation
$\text{find}~{}y\in H^{1}_{0}(D)\colon\quad A(\kappa,b,y,u)=0,\quad\text{where}\quad(\kappa,b,u)\in\mathfrak{P}_{\text{ad}}.$ (4.9)
The following result is a consequence of the implicit function theorem. The constant $C_{H^{2}}$ in the following lemma is equal to that in Lemma 4.1.
Lemma 4.6.
Let Assumption 5 hold with $\mathfrak{l}(x)\geq 0$ for a.e. $x\in D$. For every $(\kappa,b,u)\in\mathfrak{P}_{\text{ad}}$, there exists a unique $y=y(\kappa,b,u)\in H^{1}_{0}(D)$ satisfying (4.9). The mapping $\mathfrak{P}_{\text{ad}}\ni(\kappa,b,u)\mapsto y(\kappa,b,u)\in H^{1}_{0}(D)$ is infinitely many times continuously differentiable, and we have
$\|y(\kappa,b,u)\|_{H_{0}^{1}(D)}\leq C_{D}(4/\kappa_{\inf}^{*})\|b\|_{L^{2}(D)}\quad\text{for all}\quad(\kappa,b,u)\in\mathfrak{P}_{\text{ad}}.$ (4.10)
Moreover, if $u\in\mathcal{D}(\psi)$, then $y(\kappa,b,u)\in H^{1}_{0}(D)\cap H^{2}(D)$, we have
$\|y(\kappa,b,u)\|_{H^{2}(D)}\leq\frac{C_{H^{2}}\|\kappa\|_{C^{1}(\bar{D})}^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\big{(}\|\mathfrak{u}\|_{L^{\infty}(D)}C^{2}_{D}+1\big{)}\|b\|_{L^{2}(D)}.$ (4.11)
and the mapping
($\left(\left\{\,\kappa\in C^{1}(\bar{D})\;|\;\kappa>\kappa^{*}_{\inf}/2\,\right\}\times L^{2}(D)\times\mathcal{D}(\psi)\right)\ni(\kappa,b,u)\mapsto y(\kappa,b,u)\in H^{2}(D)$
is continuous.
Proof.
Given $(\kappa,b,u)\in\mathfrak{P}_{\text{ad}}$, the existence of a unique solution to (4.9) and the a priori estimate in (4.10) follow from the Lax–Milgram lemma and the definitions of $\mathfrak{P}_{\text{ad}}$ and $V_{\text{ad}}$, see also [24, sect. 7.1]. Similarly, the higher regularity and the estimate in (4.11) follow by a standard bootstrapping argument. It remains to discuss the regularity of the solution mapping. Starting with the $H^{1}_{0}(D)$-result, note that there holds
$A_{y}(\kappa,b,y,u)\delta y=A(\kappa,b,\delta y,0)\quad\text{forall}\quad\delta y\in H^{1}_{0}(D).$
Together with (4.10), this implies that $A_{y}(\kappa,b,y,u)$ is a Banach space isomorphism. Hence, the smoothness of $y(\kappa,u,b)\in H^{1}_{0}(D)$ with respect to its inputs follows from the implicit function theorem. In order to verify the $H^{2}(D)$-continuity, let
$(\kappa_{j},b_{j},u_{j})\in\left(\left\{\,\kappa\in C^{1}(\bar{D})\;|\;\kappa>\kappa^{*}_{\inf}/2\,\right\}\times L^{2}(D)\times\mathcal{D}(\psi)\right),\quad j=1,2$
denote two admissible triples and let $y_{j}\in H^{2}(D)\cap H^{1}_{0}(D)$ be the associate solutions of (4.9). The difference $\delta y=y_{1}-y_{2}\in H^{2}(D)\cap H^{1}_{0}(D)$ is the unique solution of
$\text{find}~{}y\in H^{1}_{0}(D)\colon\quad A(\kappa_{1},-A(\kappa_{1},b_{1},y_{2},u_{1}),y,u_{1})=0,$
where we identify
$A(\kappa_{1},b_{1},y_{2},u_{1})$ $=A(\kappa_{1},b_{1},y_{2},u_{1})-A(\kappa_{2},b_{2},y_{2},u_{2})$
$=-\nabla\cdot((\kappa_{1}-\kappa_{2})\nabla y_{2})+(u_{1}-u_{2})y_{2}-(b_{1}-b_{2})\in L^{2}(D).$
Invoking (4.11), we thus get
$\|y_{1}-y_{2}\|_{H^{2}(D)}\leq\frac{C_{H^{2}}\|\kappa_{1}\|_{C^{1}(\bar{D})}^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\big{(}\|\mathfrak{u}\|_{L^{\infty}(D)}C^{2}_{D}+1\big{)}\|A(\kappa_{1},b_{1},y_{2},u_{1})\|_{L^{2}(D)},$
and together with
$\|A(\kappa_{1},b_{1},y_{2},u_{1})\|_{L^{2}(D)}\leq C\left(\|\kappa_{1}-\kappa_{2}\|_{C^{1}(\bar{D})}+\|u_{1}-u_{2}\|_{L^{2}(D)}+\|b_{1}-b_{2}\|_{L^{2}(D)}\right),$
for some $C=C(y_{2})>0$, the claimed continuity result follows. ∎
From this point on, we would like to argue as in the previous section by introducing an appropriate control-to-state operator $S$ in order to define the integrand $f$. In the bilinear setting, however, this requires additional care, as the equation in (4.9) might not be well-posed for controls in a $H^{-1}(D)$-neighborhood of $B(\mathcal{D}(\psi))$. Taking this into account, we set $U_{0}=V_{\text{ad}}$ and consider the mapping
$S\colon B(U_{0})\times\Xi\to H_{0}^{1}(D),\quad\text{where}\quad S(w,\xi)=y(\kappa(\xi),b(\xi),B^{-1}_{r}w).$
Here, $B_{r}\colon U_{0}\to B(U_{0})$ is the restriction of the compact operator $B$, and $B^{-1}_{r}$ denotes is noncontinuous inverse. Using the following lemma, we deduce the Lipschitz continuity of $S$ with respect to the control.
Lemma 4.7.
If $D\subset\mathbb{R}^{d}$ is a bounded domain, then there exists a constant $C_{H^{2};H_{0}^{1}}>0$ such that for all $v\in H_{0}^{1}(D)$ and $\varphi\in H^{2}(D)$, we have $v\varphi\in H_{0}^{1}(D)$ and $\|v\varphi\|_{H_{0}^{1}(D)}\leq C_{H^{2};H_{0}^{1}}\|v\|_{H_{0}^{1}(D)}\|\varphi\|_{H^{2}(D)}.$
Proof.
This is a consequence of Theorem 7.4 in [5], the $H^{1}(D)$-trace operator’s continuity, and Friedrichs’ inequality. ∎
Lemma 4.8.
Let Assumption 5 hold with $\mathfrak{l}(x)\geq 0$ for a.e. $x\in D$. If $w_{1},w_{2}\in B(\mathcal{D}(\psi))$ and $\xi\in\Xi$, then
$\|S(w_{1},\xi)-S(w_{2},\xi)\|_{H^{1}_{0}(D)}\leq C_{H^{2};H_{0}^{1}}(4/\kappa_{\inf}^{*})\|w_{1}-w_{2}\|_{H^{-1}(D)}\|S(w_{1},\xi)\|_{H^{2}(D)}.$ (4.12)
Proof.
Let $u_{1},u_{2}\in\mathcal{D}(\psi)$ be such that $w_{j}=Bu_{j}$$j=1,2$. Since $b(\xi)\in L^{2}(D)$, we have $S(Bu_{1},\xi)\in H^{2}(D)$, see Lemma 4.6. Consequently, see Lemma 4.7$(u_{2}-u_{1})S(Bu_{1},\xi)$ induces a linear continuous functional on $H^{1}_{0}(D)$ with
$\langle(u_{2}-u_{1})S(w_{1},\xi),v\rangle_{H^{-1}(D),H_{0}^{1}(D)}$ $=\int_{D}(u_{2}-u_{1})S(u_{1},\xi)v~{}\mathrm{d}x\leq\|w_{1}-w_{2}\|_{H^{-1}(D)}\|vS(u_{1},\xi)\|_{H_{0}^{1}(D)}$
$\leq C_{H^{2};H_{0}^{1}}\|w_{1}-w_{2}\|_{H^{-1}(D)}\|S(w_{1},\xi)\|_{H^{2}(D)}\|v\|_{H_{0}^{1}(D)}$
for all $v\in H^{1}_{0}(D)$. The claimed result then follows from (4.10) noting that
$S(w_{1},\xi)-S(w_{2},\xi)=y(\kappa(\xi),(u_{2}-u_{1})S(w_{1},\xi),u_{2}).$
The following lemma verifies Assumption 1 for the integrand $f$ in (4.3).
Lemma 4.9.
Let Assumption 5 hold with $\mathfrak{l}(x)\geq 0$ for a.e. $x\in D$. Then
$0\leq f(Bu,\xi)\leq\varrho\left(C^{2}_{D}(4/\kappa_{\inf}^{*})\|b(\xi)\|_{L^{2}(D)}\right)\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi.$
Moreover, $f\colon B(U_{0})\times\Xi\to\mathbb{R}$ is continuous in its first argument on $B(\mathcal{D}(\psi))$ for each $\xi\in\Xi$ and measurable in its second argument on $\Xi$ for each $u\in B(U_{0})$.
Proof.
The measurability and continuity statements of $f$ follow from the continuity of $J$ and that of $(\kappa,b,u)\mapsto y(\kappa,b,u)$ (see Lemma 4.6), 5 (a), and the Lipschitz continuity of $S$ for fixed $\xi\in\Xi$. The stability estimate follows similarly to Lemma 4.2 by using (4.10). ∎
Next, we verify Assumptions 2 and 3 (a). We define
$C_{\text{sq}}\coloneqq\sqrt{2\ell\varrho\left(C^{2}_{D}(4/\kappa_{\inf}^{*})b_{\sup}^{*}\right)}.$
Lemma 4.10.
Let Assumption 5 hold with $\mathfrak{l}(x)\geq 0$ for a.e. $x\in D$. For each $\xi\in\Xi$, consider the mapping
$g_{\xi}\colon U_{0}\to\mathbb{R},\quad\text{where}\quad g_{\xi}(u)\coloneqq f(Bu,\xi)=J(B^{*}y(\kappa(\xi),b(\xi),u)).$
The function $g_{\xi}$ is continuous differentiable on $U_{0}$, and there holds
$\nabla g_{\xi}(u)=-p_{u,\xi}S(Bu,\xi)\in L^{2}(D),\quad\text{where}\quad p_{u,\xi}\coloneqq y(\kappa(\xi),\nabla J(B^{*}S(Bu,\xi)),u).$
Moreover, for every $(u,\xi)\in\mathcal{D}(\psi)\times\Xi$, we have $p_{u,\xi}S(Bu,\xi)\in H^{1}_{0}(D)$, the mapping
$\mathrm{D}_{w}f\colon B(\mathcal{D}(\psi))\times\Xi\to H^{1}_{0}(D),\quad\text{where}\quad\mathrm{D}_{w}f(w,\xi)\coloneqq-p_{u,\xi}S(w,\xi),$
is Carathéodory, and we have the a priori estimates
$\|\nabla g_{\xi}(u)\|_{L^{2}(D)}\leq C^{2}_{H^{1}_{0};L^{4}}C^{2}_{D}(4/\kappa_{\inf}^{*})^{2}C_{\text{sq}}\|b(\xi)\|_{L^{2}(D)}\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi,$
as well as for all $(u,\xi)\in\mathcal{D}(\psi)\times\Xi$,
$\|\mathrm{D}_{w}f(Bu,\xi)\|_{H^{1}_{0}(D)}$ $\leq C_{H^{2};H_{0}^{1}}C_{D}(4/\kappa_{\inf}^{*})\frac{C_{H^{2}}\|\kappa(\xi)\|_{C^{1}(\bar{D})}^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\big{(}\|\mathfrak{u}\|_{L^{\infty}(D)}C^{2}_{D}+1\big{)}C_{\text{sq}}\|b(\xi)\|_{L^{2}(D)}.$
Moreover, there exists an integrable random variable $L_{\mathrm{D}f}\colon\Xi\to[0,\infty)$ such that
$\|\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})\|_{L^{2}(D)}\leq L_{\mathrm{D}f}(\xi)\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}\quad\text{for all}\quad u_{1},u_{2}\in\mathcal{D}(\psi),\quad\xi\in\Xi.$
Proof.
[…]
$\|p_{u,\xi}\|_{H^{1}_{0}(D)}$ $\leq C_{D}(4/\kappa_{\inf}^{*})\|\nabla J(B^{*}S(Bu,\xi))\|_{L^{2}(D)}\leq C_{D}(4/\kappa_{\inf}^{*})\sqrt{2\ell J(B^{*}S(Bu,\xi)))}\ \leq C_{D}(4/\kappa_{\inf}^{*})C_{\text{sq}}$
for all $(u,\xi)\in U_{0}\times\Xi$, and
$\|p_{u,\xi}\|_{H^{2}(D)}\leq\frac{C_{H^{2}}\|\kappa(\xi)\|_{C^{1}(\bar{D})}^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\big{(}\|\mathfrak{u}\|_{L^{\infty}(D)}C^{2}_{D}+1\big{)}C_{\text{sq}}\quad\text{for all}\quad(u,\xi)\in\mathcal{D}(\psi)\times\Xi.$ (4.13)
As a consequence, we obtain
$\|\nabla g_{\xi}(u)\|_{L^{2}(D)}$ $\leq\|p_{u,\xi}\|_{L^{4}(D)}\|S(Bu,\xi)\|_{L^{4}(D)}\leq C^{2}_{H^{1}_{0};L^{4}}\|p_{u,\xi}\|_{H^{1}_{0}(D)}\|S(Bu,\xi)\|_{H^{1}_{0}(D)}$
$\leq C^{2}_{H^{1}_{0};L^{4}}C^{2}_{D}(4/\kappa_{\inf}^{*})^{2}C_{\text{sq}}\|b(\xi)\|_{L^{2}(D)}$
for all $(u,\xi)\in U_{0}\times\Xi$. Moreover, if $(u,\xi)\in\mathcal{D}(\psi)\times\Xi$, then
$\|\mathrm{D}_{w}f(Bu,\xi)\|_{H^{1}_{0}(D)}\leq C_{H^{2};H_{0}^{1}}\|S(Bu,\xi)\|_{H^{2}(D)}\|p_{u}\|_{H^{1}_{0}(D)}$
$\quad\leq C_{H^{2};H_{0}^{1}}C_{D}(4/\kappa_{\inf}^{*})\frac{C_{H^{2}}\|\kappa(\xi)\|_{C^{1}(\bar{D})}^{3}}{(\kappa^{*}_{\inf}/2)^{4}}\big{(}\|\mathfrak{u}\|_{L^{\infty}(D)}C^{2}_{D}+1\big{)}C_{\text{sq}}\|b(\xi)\|_{L^{2}(D)}.$
In particular, owing to Assumption 5, (4.11) and (4.13), there exists a square integrable random variable $\zeta_{\text{max}}\colon\Xi\to[0,\infty)$ such that
$\|\nabla g_{\xi}(u)\|_{L^{2}(D)}\leq\zeta_{\text{max}}(\xi)\quad\text{for all}\quad(u,\xi)\in U_{0}\times\Xi,$
as well as
$\max\big{\{}\|\mathrm{D}_{w}f(Bu,\xi)\|_{H^{1}_{0}(D)},\|S(Bu,\xi)\|_{H^{2}(D)},\|p_{u,\xi}\|_{H^{2}(D)}\big{\}}\leq\zeta_{\text{max}}(\xi)\quad\text{for all}\quad(u,\xi)\in\mathcal{D}(\psi)\times\Xi.$
In order to establish the claimed Lipschitz continuity, fix $u_{1},u_{2}\in\mathcal{D}(\psi)$ and $\xi\in\Xi$, and split
$p_{u_{1},\xi}-p_{u_{2},\xi}=p_{u_{1},\xi}-y(\kappa(\xi),\nabla J(B^{*}S(Bu_{2},\xi)),u_{1})+y(\kappa(\xi),\nabla J(B^{*}S(Bu_{2},\xi)),u_{1})-p_{u_{2},\xi}.$
For estimating the norm of the first difference, we note that
$\|\nabla J(B^{*}S(Bu_{1},\xi))-\nabla J(B^{*}S(Bu_{2},\xi))\|_{L^{2}(D)}$ $\leq\ell C_{D}\|S(Bu_{1},\xi)-S(Bu_{2},\xi)\|_{H^{1}_{0}(D)}$
$\leq\ell C_{D}C_{H^{2};H_{0}^{1}}(4/\kappa_{\inf}^{*})\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}\zeta_{\text{max}}(\xi)$
using Lemma 4.8. Similarly, arguing along the lines of proof in Lemma 4.8, we get
$\|y(\kappa(\xi),\nabla J(B^{*}S(Bu_{2},\xi)),u_{1})-p_{u_{2},\xi}\|_{H^{1}_{0}(D)}\leq C_{H^{2};H_{0}^{1}}(4/\kappa_{\inf}^{*})\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}\zeta_{\text{max}}(\xi).$
As a consequence, we have
$\|p_{u_{1}}-p_{u_{2}}\|_{H^{1}_{0}(D)}\leq C_{H^{2};H_{0}^{1}}(4/\kappa_{\inf}^{*})\zeta_{\text{max}}(\xi)\left(1+\ell C^{2}_{D}(4/\kappa_{\inf}^{*})\right)\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}.$
Finally, we decompose $\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})$ and estimate
$\|\nabla g_{\xi}(u_{1})-\nabla g_{\xi}(u_{2})\|_{H^{1}_{0}(D)}$ $\leq\left(\|p_{u_{1}}-p_{u_{2}}\|_{H^{1}_{0}(D)}+\|S(Bu_{1},\xi)-S(Bu_{2},\xi)\|_{H^{1}_{0}(D)}\right)\zeta_{\text{max}}(\xi)$
$\leq C\zeta_{\text{max}}(\xi)^{2}\|Bu_{1}-Bu_{2}\|_{H^{-1}(D)}$
for some $C>0$. By Friedrichs’ inequality, the claimed Lipschitz statement follows. ∎
Lemma 4.10 verifies Assumptions 2 and 3 (a). Moreover, as in the linear setting, 3 (b) is satisfied since $\nabla g_{\xi}$ is uniformly bounded on $\mathcal{D}(\psi)\times\Xi$.
5 Numerical illustrations
We present numerical illustrations for two instances of the risk-neutral problems analyzed in section 4. The section’s main objective is to illustrate the theoretical results established in sections 3.4 and 3.5. Before presenting numerical results, we discuss problem data, discretization aspects, implementation details, and the computation of reference solutions and gap functions.
Problem data
We chose $D=(0,1)^{2}$, the deterministic $b(x)\coloneqq 10\sin(2\pi x_{1}-x_{2})\cos(2\pi x_{2})$ in (4.2), and $J(v)\coloneqq(1/2)\|v-y_{d}\|_{L^{2}(D)}^{2}$ in (4.1) with $y_{d}(x)\coloneqq\sin(2\pi x_{1})\sin(2\pi x_{2})\exp(2x_{1})/6$. For the random diffusion coefficient $\kappa$, we used a truncated Karhunen–Loève expansion with the eigenpairs described in Example 7.56 of [17]. Our truncated expansion includes $100$ addends, uses correlation length of $1$, and independent truncated normal random variables each without shift, a standard deviation $1$, and a truncation interval of $[-3,3]$. The remaining problem data are problem-dependent and introduced in sections 5.1 and 5.2.
Reference solutions and gap functions
As the risk-neutral problems considered in section 4.1 may lack closed-form solutions, we compute reference solutions. We constructed $N_{\text{ref}}\coloneqq 8192$ samples using a fixed realization of a scrambled Sobol’ sequence, generated using scipy’s quasi-Monte Carlo engine [32]. These samples are transformed via the inverse transformation method and their subsequent values referred to reference samples. We used these samples to define a reference problem. This reference problem is an SAA problem defined by the reference samples with $N=N_{\text{ref}}$. We refer to its solutions/critical points as reference SAA solutions/critical points. Moreover, we approximate the gap functional $\Psi$ in (3.1) using the gap functional $\Psi_{\text{ref}}$ corresponding to the reference SAA problem. Similarly, we approximate the objective function $G$ via $G_{\text{ref}}$.
Discretization and implementation details
To obtain finite-dimensional optimization problems suitable for numerical solution, we discretized the state space $H_{0}^{1}(D)$ using piecewise linear continuous finite element functions and $L^{2}(D)$ using piecewise constant functions. These functions are defined on a regular triangulation of $D$ with $n\coloneqq 64$ being the number of cells in each direction. We refer to
$\min_{u\in U}\,f(Bu,\mathbb{E}[\boldsymbol{\xi}])+\psi(u)$
as the nominal problem to ($\mathcal{P}$). We solved the discretized problems using a conditional gradient method [13]. Our implementation is archived at [22]. The method stops when the gap functional falls below $10^{-10}$ or after 100 iterations, whichever comes first. Our simulation environment is built on dolfin-adjoint [7, 27], FEniCS [1, 16], and Moola [28]. The numerical simulations, except those for the nominal problems, were performed on the PACE Phoenix cluster [29]. The code and simulation output are archived at [23].
5.1 Affine-linear problem
We chose $\mathfrak{l}=-1$, $\mathfrak{u}=1$, and $\beta=0.0075$. Figure 1 depicts nominal solution and a reference SAA solution. Figure 2 depicts convergence rates for empirical estimates of the optimality gap $\mathbb{E}[G_{\text{ref}}(u_{N}^{*})-G_{\text{ref}}(u^{*})]$ and $\mathbb{E}[\|u_{N}^{*}-u^{*}\|_{L^{1}(D)}]$ over the sample size $N$. We used $40$ realizations to estimate these means, and computed convergence rates using least squares. Figure 3 depicts convergence rates of the SAA gap function’s empirical means. The empirical convergence rates closely match the theoretical ones.
Figure 1: For the affine-linear control problem, nominal solution (left), and reference SAA solution $u^{*}$ with $N=N_{\text{ref}}$ (right).
Figure 2: For the affine-linear control problem, empirical estimate of $\mathbb{E}[G_{\text{ref}}(u_{N}^{*})-G_{\text{ref}}(u^{*})]$ as a function of the sample size $N$ (left) and empirical estimate of $\mathbb{E}[\|u_{N}^{*}-u^{*}\|_{L^{1}(D)}]$ as a function of the sample size $N$ (right).
Figure 3: Empirical estimate of $\mathbb{E}[\Psi_{\text{ref}}(u_{N}^{*})]$ as a function of the sample size $N$ for the affine-linear control problem (left) and bilinear control problem (right).
5.2 Bilinear problem
We chose the constant control bounds $\mathfrak{l}=0$ and $\mathfrak{u}=1$, and $\beta=0.00055$. Figure 4 depicts nominal critical points and a reference SAA critical point. Figure 3 depicts convergence rates of the SAA gap function’s empirical means. These rates are faster than predicted by the theory, see Corollary 3.6. We think that this can be attributed to the fact that the covering number ansatz does not exploit higher-order regularity of the integrand and, in particular, potential local curvature around isolated minimizers. While a closer inspection of this improved convergence behavior is certainly of interest, it goes beyond the scope of the current work and is left for future research.
Figure 4: For the bilinear control problem, nominal critical point (left) and reference SAA critical point $u^{*}$ with $N=N_{\text{ref}}$ (right).
6 Discussion
In this paper, we have analyzed the SAA approach for risk-neutral optimization problems that incorporate a nonsmooth but convex regularization term. Our main results address the asymptotic consistency of this scheme, and the derivation of nonasymptotic sample size estimates for various optimality measures. The latter, as well as the employed techniques, come in two different flavours: for the general nonconvex case, we provide estimates on the expected value of the gap functional by applying the covering number approach. For convex objectives, and by relying on common growth conditions, we prove stronger results including convergence rates for the expected distance between minimizers, improved estimates for the gap functional, and the suboptimality in the objective function value. The presented abstract framework is applied to both linear and bilinear PDE-constrained problems under uncertainty. We also use these applications to empirically verify the sharpness of our convergence guarantees.
Our investigation also raises new questions about the SAA approach for nonsmooth minimization problems. These include, for example, the extension of the results in section 3.5 to nonconvex problems assuming suitable second-order optimality conditions and taking into account the potential existence of multiple global and/or local minimizers. Moreover, an extension of the current results to risk-averse stochastic optimization and variational inequalities may be of interest.
Reproducibility of computational results
Computer code that allows the reader to reproduce the computational results in this manuscript is available at https://doi.org/10.5281/zenodo.13336970.
Appendix A Uniform expectation bounds for expectation mappings
We establish essentially known uniform expectation bounds for integrands defined on potentially infinite-dimensional spaces. While the techniques used in this section are standard in the literature on stochastic programming, the bounds are instrumental for establishing one of our main result in section 3.4.
Let $(\Theta,\Sigma,\mu)$ be a complete probability space, and let $\boldsymbol{\zeta}$ be a random element with image space $\Theta$. Moreover, let $\boldsymbol{\zeta}^{1},\boldsymbol{\zeta}^{2},\ldots$ be sequence of independent $\Theta$-valued random elements defined on a complete probability space, each $\boldsymbol{\zeta}^{i}$ having the same distribution as $\boldsymbol{\zeta}$.
Assumption 6 (sub-Gaussian Carathéodory function, compact linear operator).
  1. (a)
    The set $V_{0}$ is a nonempty, closed, bounded, convex subset of a reflexive Banach space $V$, $W$ is a Banach space, and $B:V\to W$ is linear and compact.
  2. (b)
    The space $H$ is a separable Hilbert space, and $\mathsf{G}:B(V_{0})\times\Theta\to H$ is a Carathéodory function.
  3. (c)
    For an integrable random variable $L_{\mathsf{G}}:\Theta\to[0,\infty)$, $\mathsf{G}(\cdot,\zeta)$ is Lipschitz continuous with Lipschitz constant $L_{\mathsf{G}}(\zeta)$ for each $\zeta\in\Theta$.
  4. (d)
    There exists $\tau_{\mathsf{G}}>0$ such that for all $w\in B(V_{0})$, we have $\mathbb{E}[\exp(\tau_{\mathsf{G}}^{-2}\|\mathsf{G}(w,\boldsymbol{\zeta})-\mathbb{E}[\mathsf{G}(w,\boldsymbol{\zeta})]\|_{H}^{2})]\leq\mathrm{e}$.
We define $g(w)\coloneqq\mathbb{E}[\mathsf{G}(w,\boldsymbol{\zeta})]$, $\hat{g}_{N}(w)\coloneqq(1/N)\sum_{i=1}^{N}\mathsf{G}(w,\boldsymbol{\zeta}^{i})$, and $L\coloneqq\mathbb{E}[L_{\mathsf{G}}(\boldsymbol{\zeta})]$.
Proposition A.1.
If Assumption 6 holds, then for each $\nu>0$ and $N\in\mathbb{N}$,
$\mathbb{E}[\sup_{v\in V_{0}}\|g(Bv)-\hat{g}_{N}(Bv)\|_{H}]=\mathbb{E}[\sup_{w\in B(V_{0})}\|g(w)-\hat{g}_{N}(w)\|_{H}]\leq 2L\nu+\tfrac{\sqrt{3}\tau_{\mathsf{G}}}{\sqrt{N}}\big{(}\mathcal{N}(\nu;B(V_{0}))\big{)}^{1/2}.$
Proof.
The proof is inspired by those of Theorems 9.84 and 9.86 in [34]. We have $\|g(w_{2})-g(w_{1})\|_{H}\leq L\|w_{2}-w_{1}\|_{W}$ and $\|\hat{g}_{N}(w_{2})-\hat{g}_{N}(w_{1})\|_{H}\leq\hat{L}_{N}\|w_{2}-w_{1}\|_{W}$ for all $w_{1},w_{2}\in B(V_{0})$, where $\hat{L}_{N}\coloneqq(1/N)\sum_{i=1}^{N}L_{\mathsf{G}}(\boldsymbol{\zeta}^{i})$. We define $K\coloneqq\mathcal{N}(\nu;B(V_{0}))$. Since $B(V_{0})$ is compact, $K$ is finite and there exist $w_{1},\ldots,w_{K}\in B(V_{0})$ such that for each $w\in B(V_{0})$, we have $\|w-w_{k(w)}\|_{W}\leq\nu$, where $k(w)\coloneqq\mathrm{\arg\,\min}_{1\leq k\leq K}\,\|w-w_{k}\|_{W}$. For all $w\in B(V_{0})$,
[…]
Hence
$\sup_{v\in V_{0}}\|g(Bv)-\hat{g}_{N}(Bv)\|_{H}$ $=\sup_{w\in B(V_{0})}\|\hat{g}_{N}(w)-g(w)\|_{H}$
$\leq\hat{L}_{N}\nu+\max_{1\leq k\leq K}\,\|\hat{g}_{N}(w_{k(w)})-g(w_{k(w)})\|_{H}+L\nu.$
Using Theorem 3 in [30] and Lemma 1 in [21], we have
$\mathbb{E}[\cosh(\lambda\|\hat{g}_{N}(w_{k})-g(w_{k})\|_{H})]\leq\exp(3\lambda^{2}\tau_{\mathsf{G}}^{2}/(4N))\quad\text{for all}\quad\lambda\in\mathbb{R},\quad k=1,\ldots,K.$
Now, Lemma B.5 in [26] ensures $\mathbb{E}[\max_{1\leq k\leq K}\,\|\hat{g}_{N}(w_{k(w)})-g(w_{k(w)})\|_{H}]\leq\sqrt{3}\tau_{\mathsf{G}}\sqrt{\ln(2K)}$. Combined with $\mathbb{E}[\hat{L}_{N}]=L$, we obtain the expectation bound. ∎